a right pyramid 6 meter high has a square base of which the diagonal is root 1152 meter . volume of the pyramid is
Answer by mananth(16946)

a right pyramid 6 meter high has a square base of which the diagonal is root 1152 meter . volume of the pyramid is
Diagonal of square base sqrt(1152)
Let side of base be x m
x^2+x^2= 1152
2x^2= 1152
x^2=576
x=24
Volume of pyramid
1/3* a^2*h
1/3*24^2*6 =1152 m^3